OpenStudy (anonymous):

CHK my answers please! :) Which is the 6th term in the sequence: a_n= -1/125(5^n-1) ? -125 -25 <--- my answer -625 -5 Which is the 10th term in the sequence: -62, -48, -34, -20, -6 78 <--- my answer 50 68 64 Which represents the type of sequence: -99, -96, -92, -87, -81? Both arithmetic and geometric Neither arithmetic or geometric Geometric <--- my answer Arithmetic

OpenStudy (amistre64):

does n start at 0 or at 1?

OpenStudy (anonymous):

For which one

OpenStudy (amistre64):

for your class work ....

OpenStudy (amistre64):

and you might need to clarify the function ... is the 5^(n-1) on top or on bottom?

OpenStudy (amistre64):

also, if you give a brief discription of why you think your answers are right, it would help

OpenStudy (anonymous):

Ohh I got what u mean now! :) And the n is one the bottom and the (5^n-1) is on the top

OpenStudy (anonymous):

For my second questions aren't they all getting subtracted by. 14?

OpenStudy (amistre64):

the last 2 seem off to me

OpenStudy (amistre64):

added by 14 but yes

OpenStudy (amistre64):

-62 + 14(n-1), for n=1,2,3,4,.... n=10 gives us what?

OpenStudy (anonymous):

78?

OpenStudy (amistre64):

no, work out the grade school addition -62 + 14(9)

OpenStudy (anonymous):

I think 64 is the correct answer?

OpenStudy (amistre64):

thats better yes now the 3rd one ... define geometric and arithmetic for me

OpenStudy (anonymous):

They have to have a common difference or common ratio to be geometric or arithmetic I think

OpenStudy (amistre64):

geo is ratio, a constant multiplier add is well, addition, a constant adder is the 3rd one geo or add ?

OpenStudy (anonymous):

I think add but I'm not sure because it doesn't look like it has a common ratio either

OpenStudy (amistre64):

if we take the geoMean ... if the sqrt(a1 * a3) = a2 is a good test ... in this case:\[\sqrt{-99(-92)}\ne-96\] is -96 the arithMean of -99 and -92?

OpenStudy (amistre64):

3,4,5,6 not constant either ... so yeah, neither

OpenStudy (anonymous):

So it would be geometric because your always multiplying to find the next number in the sequence?

OpenStudy (anonymous):

Omg these confuse me so much!

OpenStudy (amistre64):

its not geometric since there is no value r such that: -99r = -96 -96r = -92

OpenStudy (anonymous):

Okay it makes a little more sense now. Thanks for the help. :)